# Absence of a bulk charge density wave signature in x-ray measurements of UTe2

### v1

@@ Methods
STM measurements have found the locations of CDW peaks on the (011) surface [3, 4, 5]. However, since STM is a surface measurement the periodicity of the CDW in the direction perpendicular to the surface is unconstrained. Using STM results we place partial constraints on the Miller indices, and identify what regions of momentum space the CDW is expected to be visible to bulk X-ray diffraction measurements, making the assumption that the wave vectors at the surface are the same as in the bulk.
@@
The peaks observed in an STM measurement [e.g., Fig. 1(b)] are projections of the bulk 3D Bragg peaks, which are indexed by Miller indices $(h,k,l)$, onto to the $(x,y)$ plane defined by Eqs. 3a and 3b. The projection is given by the two-component vector
@@
where $\textbf{q}_{3D}=\biggl{(}h\dfrac{2\pi}{a},k\dfrac{2\pi}{b},l\dfrac{2\pi}{c}\biggr{)}$. The locations of these projected points, and the associated 2D Brillouin zone of the surface, are shown in Fig. 1(c) and are consistent with STM data in Fig. 1(b). (Note that we have included only peaks with $h+k+l=2n$, where $n$ is an integer, the others being forbidden by crystal symmetry). The 2D surface-projected Brillouin zone shown in light blue in panels (b) and (c) was created by drawing perpendicular bisectors between the surface-projected Bragg peaks.
In terms of the surface basis vectors Eq. 3, the CDW modulations observed in Ref. [3], are ${q}^{\mathrm{CDW}}_{1}=\biggl{(}0.43\dfrac{2\pi}{a},\dfrac{2\pi}{\sqrt{b^{2}+c^{2}}}\biggr{)}$, ${q}^{\mathrm{CDW}}_{2}=\biggl{(}0.43\dfrac{2\pi}{a},\dfrac{-2\pi}{\sqrt{b^{2}+c^{2}}}\biggr{)}$, and ${q}^{\mathrm{CDW}}_{3}=\biggl{(}0.57\dfrac{2\pi}{a},0\biggr{)}$. The 3D positions of each of the three CDW modulations, $q_{i}^{CDW}$, such as can be constrained by a 2D STM measurement, are determined by equating,
@@
Combining Eq. 6 and 7, we find the expected locations of the three CDWs in UTe2, such as can be inferred from STM measurements, to be $\textbf{q}^{CDW}_{1}=(0.43,l^{CDW}_{1}-1,l^{CDW}_{1})$, $\textbf{q}^{CDW}_{2}=(0.43,l^{CDW}_{2}+1,l^{CDW}_{2})$, and $\textbf{q}^{CDW}_{3}=(0.57,l^{CDW}_{3},l^{CDW}_{3})$, where the values of $l^{CDW}_{i}$ are unknown. These expressions constrain the locations of the CDWs to one-dimensional cuts through 3D reciprocal space. Note that these vectors are defined relative to any crystalline Bragg peak, implying a large number of symmetry-related CDW satellites that could potentially be visible in any Brillouin zone.
Acknowledgements
@@
We thank P.F.S. Rosa for helpful discussions. This work was primarily supported by the Center for Quantum Sensing and Quantum Materials, an Energy Frontier Research Center funded by the US Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award no. DE-SC0021238. Growth of UTe2 crystals was supported by National Science Foundation grant NSF-DMR 2105191. STM measurements were supported by DOE BES award no. DE-SC0022101. Sample characterization was partly supported by DOE BES award no. DE-SC-0019154 and the Maryland Quantum Materials Center. E.F. was partly supported by NSF grant NSF-DMR 2225920. PA, VM, and JPP gratefully acknowledge additional support from the EPiQS program of the Gordon and Betty Moore Foundation, grant nos. GBMF9452, GBMF9465, and GBMF9071, respectively. This research used resources of the Advanced Photon Source, a DOE User Facility operated Argonne National Laboratory under contract no. DE-AC02-06CH11357. Identification of commercial equipment does not imply recommendation or endorsement by NIST.

### v3

@@ IV Calculation of CDW Positions
We will now focus on the regions of reciprocal space where the CDW is expected to be visible, based on what is known from the STM measurements of Refs. [25, 26, 27]. Because STM is a surface measurement, it does not constrain the periodicity of the CDW in the direction perpendicular to the surface. However, we can use the STM results to place partial constraints on the Miller indices, and identify what regions of momentum space the CDW is expected to be visible, assuming that the wave vectors at the surface are the same as in the bulk.
@@
The peaks observed in an STM measurement [e.g., Fig. 1(b)] are projections of the bulk 3D Bragg peaks, which are indexed by Miller indices $(h,k,l)$, onto to the $(x,y)$ plane defined by Eqs. 1a and 1b. The projection is given by the two-component vector
@@
where $\textbf{q}_{3D}=\biggl{(}h\dfrac{2\pi}{a},k\dfrac{2\pi}{b},l\dfrac{2\pi}{c}\biggr{)}$. The locations of these projected points, and the associated 2D Brillouin zone of the surface, are shown in Fig. 1(c) and are consistent with STM data in Fig. 1(b). (Note that we have included only peaks with $h+k+l=2n$, where $n$ is an integer, the others being forbidden by crystal symmetry). In terms of the surface basis vectors Eq. 1, the CDW modulations observed in Ref. [25], are ${q}^{\mathrm{CDW}}_{1}=\biggl{(}0.43\dfrac{2\pi}{a},\dfrac{2\pi}{\sqrt{b^{2}+c^{2}}}\biggr{)}$, ${q}^{\mathrm{CDW}}_{2}=\biggl{(}0.43\dfrac{2\pi}{a},\dfrac{-2\pi}{\sqrt{b^{2}+c^{2}}}\biggr{)}$, and ${q}^{\mathrm{CDW}}_{3}=\biggl{(}0.57\dfrac{2\pi}{a},0\biggr{)}$.
The 3D positions of each of the three CDW modulations, $q_{i}^{CDW}$, such as can be constrained by a 2D STM measurement, are determined by equating,
@@
Combining Eq. 4 and 5, we find the expected locations of the three CDW wavevectors in UTe2 derived from STM measurements to be $\textbf{q}^{CDW}_{1}=(0.43,l^{CDW}_{1}-1,l^{CDW}_{1})$, $\textbf{q}^{CDW}_{2}=(0.43,l^{CDW}_{2}+1,l^{CDW}_{2})$, and $\textbf{q}^{CDW}_{3}=(0.57,l^{CDW}_{3},l^{CDW}_{3})$, where the values of $l^{CDW}_{i}$ are unknown. These expressions constrain the locations of the CDWs to one-dimensional cuts through 3D reciprocal space. Note that these vectors are defined relative to any crystalline Bragg peak, implying a large number of symmetry-related CDW satellites that could potentially be visible in any Brillouin zone.
Note that, as far as zero-field measurements are concerned, all three of these modulations can be expressed in terms of single wave vector, $\textbf{q}^{i}_{3D}$, with the CDW being visible in different Brillouin zones with wave vector given by $\textbf{Q}_{CDW}=\textbf{G}_{3D}+\textbf{q}^{i}_{3D}$, with $\textbf{G}_{3D}$ being a reciprocal lattice vector. These different momentum-space points only become inequivalent in an applied magnetic field [25].
V Results
The wide reciprocal space cut shown in Figure 2 (c), done at fixed $h=-0.43$, should intersect the $q^{CDW}_{1}$ and $q^{CDW}_{2}$ positions from the $h=0$ zone as well as the $q^{CDW}_{3}$ position from the $h=-1$ zone. However, no CDW signal is observed above the experimental background at any of the three expected locations.
With hopes of uncovering a more subtle effect, we performed temperature difference measurements, comparing the same regions of reciprocal space at $T=1.9$ K and $T=12$ K, the latter being above the putative CDW ordering temperature [25, 26, 27]. We focused on three Brillouin zones near $(-1,1,5)$, $(-1,3,5)$, and $(-1,2,5)$, shown in Fig. 3. These data sets were integrated over a narrow range of $h$ indicated in square brackets in each panel. The red lines indicate the possible CDW locations constrained from Eqs. 4 and Eqs. 5 in each region. No changes with temperature were observed in any of these three Brillouin zones, or any of the other Brillouin zones mapped in Fig. 2.
Figure 4: Scan through (-1,2,5) Bragg peak at $T=1.9$ K in the (a) $hk$ plane and (b) $hl$ plane. Both plots are integrated ±$\pm 0.05$ r.l.u. in the out-of-plane direction. Line cut through Bragg peak along the (c) $h$, (d), $k$, and (e) $l$ directions.
Lacking a clear signature of a CDW, we can nevertheless use our data to restrict the maximum CDW amplitude in UTe2. The background level of our measurements places an upper limit on the CDW peak intensity. The majority of the background comes from the beryllium rings (clearly visible in the $hk$ plane and line cuts), whose average photon counting intensity is $\sim 10$ Hz in this reciprocal space region. Assuming the photon noise is Poisson-distributed, this places an upper bound on the maximum CDW intensity of $I_{CDW}^{max}\leq\sqrt{10}\approx 3$ Hz.
The square of the amplitude of the CDW is proportional to the integrated intensity of the CDW reflection, which can be estimated,
$I_{CDW}^{Integrated}\lesssim\frac{1}{2}(I_{CDW}^{max})(f_{CDW})^{3}$ (6)
where $f_{CDW}^{3}$ is the three-dimensional width of the CDW in reciprocal space. In the case where the CDW width is taken to be the same as in STM measurements, $f_{CDW}^{3}=(\xi^{-1}_{STM})^{3}$, where $\xi^{-1}_{STM}$ is the full-width at half-maximum of the CDW peak in STM, measured along the $q_{x}$ direction in Fig. 1(b), which we have assumed to be isotropic. In the other case, $f_{CDW}^{3}=(\delta_{h}\delta_{k}\delta_{l})$ where $\delta_{h,k,l}$ represent the widths of the nearest Bragg peak in the three orthogonal reciprocal space directions, shown in Fig. 4(c-e).
Figure 4 (a-b) shows cuts through the $(-1,2,5)$ Bragg peak. Panels (c-e) show line cuts through the maxima in the $h$, $k$, and $l$, respectively. The peak has a maximum intensity of $1.74\times 10^{6}$ Hz and widths in the $h$, $k$, and $l$ directions of $0.043\mathrm{\AA}^{-1}$, $0.032\mathrm{\AA}^{-1}$, and $0.050\mathrm{\AA}^{-1}$, respectively. The integrated intensity of this Bragg peak can then be approximated in the same way as the CDW peak shown in Eq. 6, $I_{Bragg}^{Integrated}=\frac{1}{2}(I_{Bragg}^{max})(\delta_{h}\delta_{k}\delta_{l})$, where $\delta_{h,k,l}$ are the measured widths along the crystallographic directions. Using this number, we can place an upper limit on the amplitude charge density of the CDW,
where $\rho_{(-1,2,5)}$ is the calculated structure factor of the $(-1,2,5)$ Bragg peak and $\rho_{CDW}$ is the charge amplitude of the CDW. The structure factor of the $(-1,2,5)$ Bragg peak is calculated to be $\rho_{(-1,2,5)}=169.38\>\mathrm{e}^{-}/v_{\mathrm{cell}}$, where $v_{\mathrm{cell}}$ is the volume of the standard orthorhombic unit cell. Comparing this to the amplitude of the charge density of the CDW gives $\rho_{CDW}\lesssim 2.13\>e^{-}/v_{\mathrm{cell}}$ in the case that the CDWs have the same width as in STM measurements or $\rho_{CDW}\lesssim 0.70\>e^{-}/v_{\mathrm{cell}}$ in the case that the CDWs have the same width as the (-1,2,5) Bragg peak.
VI Conclusions
Our study indicates that UTe2 does not exhibit a bulk CDW of the sort that occurs in many other transition-metal dichalcogenides, such as NbSe2 or TaS2 [38], which exhibit pronounced CDW signatures in both STM and x-ray measurements [32, 39]. Note that, while we have used STM measurements as a guide for where to look in reciprocal space, we performed a 3D momentum sweep covering $>20$ Brillouin zones, and we would have observed the CDW even if its wave vector deviated from what was observed in STM.
How to interpret the upper bound on the amplitude, $\rho_{CDW}$, depends on whether one assumes the CDW is electronic (the scattering arising purely from a valence modulation) or structural (the scattering arising from a periodic distortion to the crystalline lattice).
Assuming the CDW is purely structural, the magnitude of the lattice displacement should be of order $u_{max}/a\sim\rho_{CDW}/\rho_{tot}$, where $\rho_{tot}$ represents the total charge density [38]. Assuming the order is long-ranged (see Section V), the amplitude bound $\rho_{CDW}\lesssim 0.70\>e^{-}/v_{\mathrm{cell}}$ translates into an atomic displacement $u_{max}/a\lesssim 8.93\times 10^{-4}$, where $a$ is the lattice parameter of UTe2. Alternatively, if we now consider the case in which the CDW has a short correlation length, the atomic displacement would be $u_{max}/a\lesssim 2.72\times 10^{-3}$. Both of these numbers are markedly lower than the atomic displacements in other dichalcogenide CDWs, which have distortions $u_{max}/a$ ranging from 1-7% [38]. We conclude that whatever CDW was observed in STM experiments in UTe2 has a completely different character from CDWs in other dichalcogenides.
We can also consider the case in which the CDW is assumed to be purely electronic, meaning the atoms do not shift from their equilibrium positions, and the CDW is dominated by a modulation in the valence electron density [34, 40, 41]. In this case, our x-ray measurements allow for an amplitude of up to $\sim 2\>e^{-}$ per unit cell. This would represent a significant fraction of the valence band [13, 42, 43]. Such an amplitude might be comparable to the effect seen in STM, though it would be difficult to reconcile with the absence of a resistive transition in bulk transport measurements in this range of temperatures [11, 44, 45]. We cannot, however, completely rule out this possibility.
A final explanation is that the CDW observed in STM measurements is restricted to the surface of UTe2. This would explain why its presence is so robust in multiple STM measurements [25, 26, 27] and absent in transport measurements and the bulk-sensitive x-ray measurements presented here. This scenario would be reminiscent of the CDW seen in STM measurements of NaxCa2-xCuO2Cl2 [46]. Such a CDW transition can be explained through the framework of extraordinary phase transitions, in which the critical behavior is different at the surface compared to in the bulk of material [47, 48]. As a result the transition temperature of the surface can be higher than that of the bulk. This allows for the possibility that the CDW in UTe2 could be of either electronic or structural origin at the surface, with the absence of a bulk signature in x-ray and transport measurements being a consequence of the small volume fraction of the effect. In this scenario, one might expect different surface terminations to lead to different surface ordering effects.
We would like to note that two other reports of experimental investigations of a resonant hard x-ray diffraction experiment [49] and of ultrasound measurements [50] of UTe2 at low temperature, were reported simultaneously with this work. All measurements, performed independently and on samples grown by more than one group, arrive at the same conclusions reported here.
VII Acknowledgements
@@
We thank Priscila F.S. Rosa for helpful discussions. CSK acknowledges useful discussions with Julian May-Mann. This work was primarily supported by the Center for Quantum Sensing and Quantum Materials, an Energy Frontier Research Center funded by the US Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award no. DE-SC0021238. Growth of UTe2 crystals was supported by National Science Foundation grant NSF-DMR 2105191. STM measurements were supported by DOE BES award no. DE-SC0022101. Sample characterization was partly supported by DOE BES award no. DE-SC-0019154 and the Maryland Quantum Materials Center. E.F. was partly supported by NSF grant NSF-DMR 2225920. PA, VM, and JPP gratefully acknowledge additional support from the EPiQS program of the Gordon and Betty Moore Foundation, grant nos. GBMF9452, GBMF9465, and GBMF9071, respectively. This research used resources of the Advanced Photon Source, a DOE User Facility operated Argonne National Laboratory under contract no. DE-AC02-06CH11357. Identification of commercial equipment does not imply recommendation or endorsement by NIST.